OpenStudy (anonymous):

Solve the linear system. 15x + 10y = 145 and 4x - 6y = 56

OpenStudy (johnweldon1993):

You can use either the elimination method, or the substitution method here to solve these...

OpenStudy (jhannybean):

Solve 1 equation for y first. \[\large 4x-6y=56 \to y= \frac{ -4x+56 }{ -6 } = \frac{ 2 }{ 3 }x-\frac{ 28 }{ 3 }\]\] use this to plug into the other equation to solve for one variable.\[\large 15x+10(\frac{ 2 }{ 3 }x-\frac{ 28 }{ 3 })=145\]\[\large 15x+\frac{ 20 }{ 3 }x-\frac{ 280 }{ 3 }=145\]\[\large \frac{ 65 }{ 3 }x=\frac{ 715 }{ 3 }\]\[x=\frac{ 715 }{ 3 }*\frac{ 3 }{ 65}=11\] So we found x, then plug "x=11" back into any equation to solve for y, you will have both of your variables :)

OpenStudy (anonymous):

solution: x = 11, y= -2 -------------------------------------------------- 150 x y = 145

OpenStudy (jhannybean):

ok? I wasn't giving solutions, just trying to help the person work out the problem, but thanks.

OpenStudy (anonymous):

Oh sorry haha

OpenStudy (jhannybean):

Alrighty.

OpenStudy (anonymous):

thanks both of you

OpenStudy (anonymous):

i like the solution part tho haha

OpenStudy (jhannybean):

You mean the answers.

OpenStudy (anonymous):

No problemo haha